### v1

@@ Abstract
The rising popularity of multimodal large language models (MLLMs) has sparked a significant increase in research dedicated to evaluating these models. However, current evaluation studies predominantly concentrate on the ability of models to comprehend and reason within a unimodal (vision-only) context, overlooking critical performance evaluations in complex multimodal reasoning tasks that integrate both visual and text contexts. Furthermore, tasks that demand reasoning across multiple modalities pose greater challenges and require a deep understanding of multimodal contexts. In this paper, we introduce a comprehensive assessment framework named MM-InstructEval, which integrates a diverse array of metrics to provide an extensive evaluation of the performance of various models and instructions across a broad range of multimodal reasoning tasks with vision-text contexts. MM-InstructEval enhances the research on the performance of MLLMs in complex multimodal reasoning tasks, facilitating a more thorough and holistic zero-shot evaluation of MLLMs. We firstly utilize the “Best Performance” metric to determine the upper performance limit of each model across various datasets. The “Mean Relative Gain” metric provides an analysis of the overall performance across different models and instructions, while the “Stability” metric evaluates their sensitivity to variations. Historically, the research has focused on evaluating models independently or solely assessing instructions, overlooking the interplay between models and instructions. To address this gap, we introduce the “Adaptability” metric, designed to quantify the degree of adaptability between models and instructions. Evaluations are conducted on 31 models (23 MLLMs) across 16 multimodal datasets, covering 6 tasks, with 10 distinct instructions. The extensive analysis enables us to derive novel insights.
@@
Multimodal Large Language Models Multimodal Reasoning Tasks with Vision-Text Contexts Various Instructions Zero-shot Evaluation Comprehensive Metrics
@@
1 Introduction
@@
Multimodal Large Language Models (MLLMs), harnessing the formidable capabilities of Large Language Models (LLMs), demonstrate outstanding performance across a spectrum of multimodal tasks (Alayrac et al., 2022; Gao et al., 2023; Li et al., 2023b; Dai et al., 2023). Recent research developments, as summarized in Table 1, include but are not limited to MME (Fu et al., 2023), MMBench (Liu et al., 2023c), SEED-Bench (Li et al., 2023a), LVLM-eHub (Xu et al., 2023), MM-Vet (Yu et al., 2023b), MMMU (Yue et al., 2023), and MLLM-Bench (Ge et al., 2023). These studies primarily focus on evaluating the traditional vision-language multimodal capabilities of MLLMs in tasks predominantly driven by visual context (vision context + text question), such as Visual Question Answering (VQA) and Video Question Answering (VideoQA). In these tasks, the answers to the questions are often derived solely based on the image context, as illustrated in the section above the dotted line in Figure 1. These capabilities include Recognition, Optical Character Recognition (OCR), Spatial Reasoning, Knowledge Integration, and Code Reasoning (Yu et al., 2023b). However, the performance of Multimodal Large Language Models (MLLMs) in multimodal reasoning tasks that combine vision-text contexts (vision-text contexts + text question) remains less explored. Such tasks include Visual Question Answering with Multimodal Contexts (VQAMC) (Lu et al., 2022a; Yue et al., 2023; Ghosal et al., 2024), Multimodal Sentiment Analysis (MSA) (Niu et al., 2016; Yang et al., 2021a; Zadeh et al., 2016, 2018), Multimodal Aspect-Based Sentiment Analysis (MABSA) (Zhang et al., 2018; Lu et al., 2018; Zhou et al., 2021), Multimodal Hateful Memes Detection (MHMD) (Mathias et al., 2021), Multimodal Sarcasm Detection (MSD) (Cai et al., 2019), and Multimodal Relation Extraction (MRE) (Mathias et al., 2021). In the field of Natural Language Processing (NLP), most research (Zhang et al., 2023b; Wang et al., 2023; Dalvi et al., 2023) primarily evaluates pure Large Language Models (LLMs) like ChatGPT (OpenAI, 2023a), Flan-T5 (Chung et al., 2022), focusing specifically on text classification tasks such as text sentiment analysis and relation classification. It leaves a gap in understanding the performance of various MLLMs in multimodal reasoning tasks that rely on both text and image contexts. For example, in the Visual Question Answering with Multimodal Contexts (VQAMC) task depicted in Figure 1, it is clear that the answer to the question depends on integrating the image-text contexts. Simply using the image and the question without incorporating the textual context fails to produce the correct answer.
To address the aforementioned gap, we conduct the comprehensive evaluation involving 31 publicly available models, including 23 Multimodal Large Language Models (MLLMs), across a diverse set of 16 datasets covering 6 distinct tasks, as detailed in Tables 2 and 3. Our primary objective is to assess the performance of various MLLMs for multimodal reasoning tasks that integrate vision-text contexts. Additionally, we aim to establish benchmarks for a range of MLLMs across diverse multimodal reasoning tasks. These tasks demand not only traditional vision-language multimodal capabilities from the models but also require a deep understanding of multimodal contexts for effective reasoning to arrive at the final answer. It is illustrated in the section below the dotted line in Figure 1. Multimodal reasoning tasks with vision-text contexts often necessitate various cognitive processes, including but not limited to multimodal alignment, multimodal interaction, and multimodal fusion, as emphasized in studies (Han et al., 2021; Hazarika et al., 2020; Wu and Li, 2023), alongside content understanding (Wu et al., 2024). Defining the boundaries of such multimodal reasoning with vision-text contexts can be challenging, as the multimodal contexts may encompass diverse elements such as visual-language alignment through OCR and spatial reasoning, among others. Furthermore, in tasks like visual question answering with multimodal contexts, especially when accompanied by prompts or instructions, these reasoning abilities are often triggered, as vividly demonstrated in datasets like AlgoPuzzleVQA (Ghosal et al., 2024).
@@
Figure 2: Overview of our MM-InstructEval framework, which conducts evaluations of popular MLLMs across various multimodal reasoning tasks with multimodal contexts utilizing comprehensive metrics. As illustrated in the dashed box, we select the ‘MVSA-Single’ dataset from the ‘MSC’ task to utilize ‘Instruction # 2’ for evaluating a specific ‘MLLM’. We then aggregate results to thoroughly assess the performance of models and instructions using a variety of metrics. A colorful ‘Q’ symbolizes ‘Question’, and its design varies according to the specific task. For more detailed visual representations and explanations, please refer to Figure 5.
We introduce a comprehensive assessment framework named MM-InstructEval, which incorporates a diverse set of metrics to thoroughly evaluate various models and instructions within the realm of multimodal reasoning tasks with vison-text contexts, as depicted in Figure 2. MM-InstructEval is designed to complement existing zero-shot evaluation studies of Multimodal Large Language Models (MLLMs), offering a more comprehensive and holistic assessment in conjunction with prior research. Specifically, we introduce the “Best Performance” metric to explore how each model achieves its peak performance under specific instructions for each dataset. The metric establishes the upper benchmark for capabilities of each model on a given dataset. To assess the overall performance of each model across all instructions on a particular dataset, we propose the “Model Mean Relative Gain” metric. Similarly, the “Instruction Mean Relative Gain” metric is designed to evaluate the effectiveness of each instruction across all models. Stability is also a critical indicator of both model and instruction efficacy. A model that consistently delivers strong performance across various instructions is deemed more stable, while an instruction that consistently yields high performance across different models is considered highly effective. To measure robustness of models and instructions possessing superior performance, we introduce the “Model Stability” and “Instruction Stability” metrics. Furthermore, prior studies have typically focused either solely on assessing different models (Fu et al., 2023; Li et al., 2023a; Xu et al., 2023; Yu et al., 2023b; Yue et al., 2023; Ge et al., 2023) or exclusively on evaluating the performance of instructions (Chia et al., 2023; Ajith et al., 2023), often neglecting the crucial aspect of adaptability between models and instructions. To address this gap, we introduce the “Adaptability” metric to quantify how well different models adapt to various instructions. The metric evaluates how frequently each instruction achieves Top-K performance on a specific model across all datasets. Thereby, we can pinpoint which instructions consistently deliver superior performance for each model, effectively bridging the evaluation gap. Adaptability metric provides key insights into how models adjust to different instructional contexts, thereby enhancing the utility of MLLMs in practical applications. We conduct extensive experiments, and the results reveal significant insights into the performance of large (multimodal) language models in multimodal reasoning tasks with vision-text contexts: (1) Open-source vs. Closed-source: In most tasks, the performance of advanced closed-source models, such as GPT-4V and Claude3, generally exceeds that of all open-source models. Notably, these closed-source models demonstrate significantly superior performance in more challenging scenarios, such as the multimodal relation extraction task, etc. (2) Model Architecture and Performance: In the open-source group, models utilizing the Flan-T5 backbone (Chung et al., 2022), which is based on the Encoder-Decoder architecture, such as BLIP-2 (Li et al., 2023b) and InstructBLIP (Dai et al., 2023), outperform those in the LLaMA series (Touvron et al., 2023a, b), which employ a Decoder-only architecture. It suggests that the Encoder-Decoder framework may be more adept at handling the multimodal reasoning tasks possessing vison-text contexts. (3) Effectiveness of Instruction Formats: Instructions formatted as “Question-Answer” consistently yield better results. Furthermore, the inclusion of options within these instructions enhances performance on specific tasks, highlighting the importance of structured interaction in improving task comprehension and response accuracy. (4) Instruction Preferences to Models: Due to varied training paradigms, different models demonstrate distinct preferences for specific instructions. The variability underscores the need for adaptable and flexible instruction strategies to maximize model performance on diverse tasks. Our main contributions are summarized as follows:
  • We conduct zero-shot evaluations on 31 models across 16 datasets, utilizing 10 distinct instructions to cover 6 diverse multimodal reasoning tasks that depend on vision-text contexts. The comprehensive assessment, the first of its kind, evaluates a broad spectrum of MLLMs across various instructions in complex multimodal reasoning tasks. Our work builds upon and significantly extends previous evaluations of MLLMs, culminating in a thorough appraisal of cutting-edge MLLMs.
@@
  • We introduce the MM-InstructEval framework, which incorporates a variety of metrics to provide a detailed evaluation of different models and instructions. Comprehensive metrics include the Best Performance metric to benchmark peak model capabilities, the Mean Relative Gain metric to assess overall efficacy, and the Stability metric to measure the robustness. Additionally, the Adaptability metric is proposed to quantify the adaptability between models and instructions, offering valuable insights for the design of instructions to maximize the performance of current and future models.
  • Through extensive experimentation, we establish benchmarks for both LLMs and MLLMs across various multimodal reasoning tasks with vision-text contexts. Our findings enrich the understanding of the interactions between model architectures, instruction formats, and task performance, providing strategic guidance for future advancements in applications of MLLMs.
@@ 3.2 Various Multimodal Instructions
Recent research in Natural Language Processing (NLP) has demonstrated that variations in instructions, even with identical semantics, can significantly affect the performance of a model (Perez et al., 2021; Lu et al., 2022b; Zhang et al., 2023b; Liu et al., 2023b; Gan and Mori, 2023). Building on the insights from Zhang et al., we craft multimodal instructions for diverse multimodal reasoning tasks incorporating vision-text contexts, which include essential elements such as the task name, definition, and output format, as well as optional components like answer options. The text and image contexts form the multimodal content for each instance across various multimodal reasoning tasks. These instructions are specifically designed to evaluate how different MLLMs perform in a zero-shot setting when faced with diverse instructions. Although we develop only 10 manual instructions to gauge the performance of various models across different datasets, the evaluations offer valuable insights, enabling us to draw significant conclusions. For instance, we design a range of multimodal instructions for the Visual Question Answering with Multimodal Contexts task, as depicted in Figure 3. Similarly, consistent instruction structures that integrate both text and image contexts also are created for other tasks. The instructions tailored for the Multimodal Sarcasm Detection (MSD) task are detailed in Figure 4.
@@ 3.3 Comprehensive Metrics
$p^{j}_{mdi}=LM(T^{j},V^{j}),\ acc_{mdi}=\frac{\sum_{j=1}^{N_{d}}{\mathbbm{1}}_{(p^{j}_{mdi}=l^{j}_{mdi})}}{N_{d}},$ (1)
where LM represents a specific language model, the LLM takes only the text instruction, including text context, as input, denoted as $T$, while the MLLM takes the multimodal context, $(T,V)$, as input. For the $j$-th instance, $p^{j}_{mdi}$ represents the predicted label of the LM, $l^{j}_{mdi}$ is the true label. $\mathbbm{1}$ means the indicator function, $N_{d}$ is the number of instances, $acc_{mdi}\in Acc_{mdi}$ is the accuracy performance, $Acc_{mdi}\in\mathbb{R}^{|\mathcal{M}|\times|\mathcal{D}|\times|\mathcal{I}|}$, |||*|| ∗ | means the number of *. $|\mathcal{M}|=31$, $|\mathcal{D}|=16$, and $|\mathcal{I}|=10$ represent the number of evaluated models, datasets, and instructions, respectively.
@@ 4.2 Evaluated Models
Table 2: Comprehensive summary of various models (Red indicates closed-source models; others are open-source). Below are the abbreviations used in the table: ‘LLMs’ refers to the pretrained Language Models (LLMs) that form the backbone of the Multimodal Large Language Models (MLLMs). ‘PVM’ signifies the pretrained visual model backbone of the MLLMs. ‘To-Paras’ and ‘Tr-Paras’ represent the total number of parameters and trainable parameters for each language model, respectively. ‘Held-In’ refers to the dataset that each specific MLLM was trained or fine-tuned on. The ‘GPU’ column indicates the single GPU utilization during inference, and the ‘Time’ column signifies the time taken for model inference on each text/multimodal instance using a single GPU. Note that GPU usage and inference time may vary slightly across different datasets due to varying data lengths. For these two metrics, we provide an approximate mean value across all datasets for each model. A dash, ‘-’, indicates that the data is not applicable or not involved, and ‘Unknown’ means we do not have information on whether these models cover the evaluated dataset.
@@
Multimodal GPT-4V gpt-4 vision-preview - - Unknown - -
@@
Claude3 claude-3-opus-$2$ Unknown - -
OpenFlamingo LLaMA-7B ViT-L/14 8.34B 1.31B 34G 1.5s
@@
Multimodal BLIP-2 Flan-T5-XXL ViT-G/14 12.23B 0.11B - 26G 3.5s
InstructBLIP Flan-T5-XXL ViT-G/14 12.31B 0.45B - 16G 0.3s

### v2

@@ Abstract
The emergence of multimodal large language models (MLLMs) has triggered extensive research in model evaluation. While existing evaluation studies primarily focus on unimodal (vision-only) comprehension and reasoning capabilities, they overlook critical assessments of complex multimodal reasoning tasks that require integrated understanding of both visual and textual contexts. Such multimodal tasks present unique challenges, demanding sophisticated reasoning across multiple modalities and deep comprehension of multimodal contexts. In this paper, we present MM-InstructEval, a comprehensive evaluation framework that incorporates diverse metrics to assess model performance across various multimodal reasoning tasks with vision-text contexts. We conduct extensive zero-shot evaluations on 45 models (including 36 MLLMs) across 16 multimodal datasets, encompassing 6 distinct tasks using 10 different instructions. Our framework introduces multiple innovative metrics, including the “Best Performance” metric to benchmark peak model capabilities, the “Mean Relative Gain” metric to assess overall efficacy across models and instructions, the “Stability” metric to measure robustness, and the “Adaptability” metric to quantify the compatibility between models and instructions. Through comprehensive evaluation and analysis, we uncover several significant insights about model architectures, instruction formats, and their interactions in multimodal reasoning tasks. Our findings establish new benchmarks for assessing the reasoning capabilities of MLLMs and provide strategic guidance for future developments. To facilitate continued research and evaluation in this field, we release our framework and resources at https://github.com/declare-lab/MM-InstructEval, with an interactive leaderboard available at MM-InstructEval Leaderboard.
@@
[…]
1 Introduction
@@
Multimodal Large Language Models (MLLMs), harnessing the formidable capabilities of Large Language Models (LLMs), demonstrate outstanding performance across a spectrum of multimodal tasks (Alayrac et al., 2022; Gao et al., 2023; Li et al., 2023b; Dai et al., 2023). Recent research developments, as summarized in Table 1, include but are not limited to MME (Fu et al., 2023), MMBench (Liu et al., 2023c), SEED-Bench (Li et al., 2023a), LVLM-eHub (Xu et al., 2023), MM-Vet (Yu et al., 2023b), MMMU (Yue et al., 2023), and MLLM-Bench (Ge et al., 2023). These studies primarily focus on evaluating the traditional vision-language multimodal capabilities of MLLMs in tasks predominantly driven by visual context (vision context + text question), such as Visual Question Answering (VQA) and Video Question Answering (VideoQA). In these tasks, the answers to the questions are often derived solely based on the image context, as illustrated in the section above the dotted line in Figure 1. These capabilities include Recognition, Optical Character Recognition (OCR), Spatial Reasoning, Knowledge Integration, and Code Reasoning (Yu et al., 2023b). However, the performance of Multimodal Large Language Models (MLLMs) in multimodal reasoning tasks that combine vision-text contexts (vision-text contexts + text question) remains less explored. Such tasks include Visual Question Answering with Multimodal Contexts (VQAMC) (Lu et al., 2022a; Yue et al., 2023; Chia et al., 2024), Multimodal Sentiment Analysis (MSA) (Niu et al., 2016; Yang et al., 2021a; Zadeh et al., 2016, 2018), Multimodal Aspect-Based Sentiment Analysis (MABSA) (Zhang et al., 2018; Lu et al., 2018; Zhou et al., 2021), Multimodal Hateful Memes Detection (MHMD) (Mathias et al., 2021), Multimodal Sarcasm Detection (MSD) (Cai et al., 2019), and Multimodal Relation Extraction (MRE) (Mathias et al., 2021). In the field of Natural Language Processing (NLP), most research (Zhang et al., 2023b; Wang et al., 2023; Dalvi et al., 2023) primarily evaluates pure Large Language Models (LLMs) like ChatGPT (OpenAI, 2023a), Flan-T5 (Chung et al., 2022), focusing specifically on text classification tasks such as text sentiment analysis and relation classification. It leaves a gap in understanding the performance of various MLLMs in multimodal reasoning tasks that rely on both text and image contexts. For example, in the Visual Question Answering with Multimodal Contexts (VQAMC) task depicted in Figure 1, it is clear that the answer to the question depends on integrating the image-text contexts and multi-step reasoning. Simply using the image and the question without incorporating the textual context fails to produce the correct answer.
To bridge this research gap, we present a comprehensive evaluation of 45 publicly available models, including 36 MLLMs, across 16 datasets encompassing 6 distinct tasks. These tasks demand not only traditional vision-language capabilities but also sophisticated reasoning abilities to integrate and understand multimodal contexts effectively. The complexity of such tasks is exemplified in VQAMC, where correct responses require the integration of both visual and textual information, as simple image-question pairs without textual context prove insufficient. Our evaluation framework addresses the multifaceted nature of multimodal reasoning with vision-text contexts, which encompasses various cognitive processes, including multimodal alignment, interaction, and fusion (Han et al., 2021; Hazarika et al., 2020; Wu & Li, 2023), alongside content understanding (Wu et al., 2024a). While defining the boundaries of such reasoning can be challenging due to the diverse elements involved (e.g., visual-language alignment, OCR, spatial reasoning), datasets like AlgoPuzzleVQA (Chia et al., 2024) demonstrate how these reasoning capabilities are activated through carefully designed prompts and instructions.
@@
We present MM-InstructEval, a comprehensive assessment framework that integrates diverse metrics to evaluate models and instructions in multimodal reasoning tasks with vision-text contexts (Figure 2). This framework complements existing zero-shot evaluations of MLLMs, offering a more holistic assessment approach through several innovative metrics: The “Best Performance” metric establishes the peak capabilities of each model under optimal instructions for specific datasets. To capture overall effectiveness, we introduce the “Model Mean Relative Gain” metric for assessing model performance across instructions, and the “Instruction Mean Relative Gain” metric for evaluating instruction effectiveness across models. The “Model Stability” and “Instruction Stability” metrics measure robustness by evaluating consistency of performance across different scenarios. While previous studies have focused separately on either model evaluation (Fu et al., 2023; Li et al., 2023a; Xu et al., 2023; Yu et al., 2023b; Yue et al., 2023; Ge et al., 2023) or instruction assessment (Chia et al., 2023; Ajith et al., 2023), we propose the “Adaptability” metric to quantify how well different models adapt to various instructions. This metric measures how frequently instructions achieve Top-K performance across datasets for specific models, providing crucial insights into optimal instruction design for different models.
Our extensive experiments reveal significant insights into the performance of large (multimodal) language models in multimodal reasoning tasks with vision-text contexts: First, advanced closed-source models (GPT-4V, GLM-4V-plus) consistently outperform open-source alternatives across most tasks, with particularly notable performance gaps emerging in challenging scenarios like multimodal relation extraction and the VQAMA task requiring multi-step reasoning. Within open-source models, Encoder-Decoder architectures (using Flan-T5 backbone, e.g., BLIP-2, InstructBLIP) demonstrate superior performance compared to the LLaMA series. Second, the latest models, such as Qwen2.5-VL, MiciCPM-V2.6, and others, showcase exceptional performance through architectural innovations and improved training strategies. Notably, newer models achieve comparable results with significantly fewer parameters (e.g., Qwen2.5-VL-7B-Instruct vs. Qwen2-VL-72B-Instruct), indicating higher information density and more effective parameter utilization in latest models. Third, regarding instruction format impact, “Question-Answer” format consistently yields superior results, and including structured options within instructions further enhances performance on specific tasks, providing practical guidance for optimizing model interactions in real-world applications. Fourth, different models show varying preferences for instruction formats, reflecting their diverse training paradigms, which emphasizes the importance of flexible instruction strategies for optimal performance and is crucial for maximizing model effectiveness across different tasks. Fifth, while current open-source MLLMs excel in basic vision-text reasoning, they struggle with more complex scenarios, showing significant performance gaps in multi-step reasoning tasks (e.g., PuzzleVQA), fine-grained information extraction, and long-tail phenomena in multimodal relation extraction. These findings provide valuable insights for both theoretical understanding and practical applications of MLLMs, while also identifying key areas for future development in complex reasoning capabilities and model architecture design.
@@
Figure 2: Overview of our MM-InstructEval framework, which conducts evaluations of popular MLLMs across various multimodal reasoning tasks with multimodal contexts utilizing comprehensive metrics. As illustrated in the dashed box, we select the ‘MVSA-Single’ dataset from the ‘MSC’ task to utilize ‘Instruction # 2’ for evaluating a specific ‘MLLM’. We then aggregate results to thoroughly assess the performance of models and instructions using a variety of metrics. A colorful ‘Q’ symbolizes ‘Question, ’ and its design varies according to the specific task. For more detailed visual representations and explanations, please refer to Figure 5.
Our main contributions are summarized as follows:
@@
  • We introduce the MM-InstructEval framework, which incorporates multiple metrics for detailed model and instruction evaluation, including the Best Performance metric for benchmarking peak capabilities, the Mean Relative Gain metric for assessing overall efficacy, the Stability metric for measuring robustness, and the Adaptability metric for quantifying model-instruction compatibility, offering valuable insights for instruction design to maximize current and future model performance.
  • We conduct comprehensive zero-shot evaluations of 45 models across 16 datasets, employing 10 distinct instructions to assess 6 diverse multimodal reasoning tasks that require integrated vision-text understanding. This represents the first extensive assessment of MLLMs across various instructions, with a particular focus on complex multimodal reasoning scenarios.
  • Through extensive experimentation, we establish comprehensive benchmarks for both LLMs and MLLMs across various multimodal reasoning tasks with vision-text contexts. We significantly expand upon previous MLLM evaluations by introducing systematic comparisons across models, datasets, and instructions. This extensive evaluation uncovers critical insights into model behaviors, limitations, and performance variability, providing practitioners with practical guidance for model selection and instruction design to optimize performance in multimodal tasks.
@@ 3.2 Various Multimodal Instructions
Recent research in Natural Language Processing (NLP) has demonstrated that variations in instructions, even with identical semantics, can significantly affect the performance of a model (Perez et al., 2021; Lu et al., 2022b; Zhang et al., 2023b; Liu et al., 2023b; Gan & Mori, 2023). Building on the insights from Zhang et al., we craft multimodal instructions for diverse multimodal reasoning tasks incorporating vision-text contexts, which include essential elements such as the task name, definition, and output format, as well as optional components like answer options. The text and image contexts form the multimodal content for each instance across various multimodal reasoning tasks. These instructions are specifically designed to evaluate how different MLLMs perform in a zero-shot setting when faced with diverse instructions. Although we develop only 10 manual instructions to gauge the performance of various models across different datasets, the evaluations offer valuable insights, enabling us to draw significant conclusions. For instance, we design a range of multimodal instructions for the Visual Question Answering with Multimodal Contexts task, as depicted in Figure 3. Similarly, consistent instruction structures that integrate both text and image contexts also are created for other tasks. The instructions tailored for the Multimodal Sarcasm Detection (MSD) task are detailed in Figure 4.
@@ 3.3 Comprehensive Metrics
$p^{j}_{mdi}=LM(T^{j},V^{j}),\ acc_{mdi}=\frac{\sum_{j=1}^{N_{d}}{\mathbb{1}}_{(p^{j}_{mdi}=l^{j}_{mdi})}}{N_{d}},$ (1)
where LM represents a specific language model, the LLM takes only the text instruction, including text context, as input, denoted as $T$, while the MLLM takes the multimodal context, $(T,V)$, as input. For the $j$-th instance, $p^{j}_{mdi}$ represents the predicted label of the LM, $l^{j}_{mdi}$ is the true label. $\mathbb{1}$ means the indicator function, $N_{d}$ is the number of instances, $acc_{mdi}\in Acc_{mdi}$ is the accuracy performance, $Acc_{mdi}\in\mathbb{R}^{|\mathcal{M}|\times|\mathcal{D}|\times|\mathcal{I}|}$, |||*|| ∗ | means the number of *. $|\mathcal{M}|=35$, $|\mathcal{D}|=16$, and $|\mathcal{I}|=10$ represent the number of evaluated models, datasets, and instructions, respectively.
@@ 4.2 Evaluated Models
Table 2: A comprehensive summary of various models (Red indicates closed-source models; others are open-source). Below are the abbreviations used in the table: ‘LLMs’ refers to the pretrained Language Models (LLMs) that form the backbone of the Multimodal Large Language Models (MLLMs). ‘PVM’ signifies the pretrained visual model backbone of the MLLMs. ‘To-Paras’ and ‘Tr-Paras’ represent the total number of parameters and trainable parameters for each language model, respectively. ‘Held-In’ refers to the dataset that each specific MLLM was trained or fine-tuned on. The ‘GPU’ column indicates the single GPU utilization during inference, and the ‘Time’ column signifies the time taken for model inference on each text/multimodal instance using a single GPU. Note that GPU usage and inference time may vary slightly across different datasets due to varying data lengths. For these two metrics, we provide an approximate mean value across all datasets for each model. A dash, ‘-’, indicates that the data is not applicable or not involved, and ‘Unknown’ means we do not have information on whether these models cover the evaluated dataset.
@@
[…]
GPT-4$\heartsuit$ gpt-4o - - - Unknown - -
Claude3$\heartsuit$ claude-3-opus-$2$ Unknown - -
GLM-4V-$\heartsuit$ GLM-4 - - - Unknown - -
@@
Open MLLMs OpenFlamingo LLaMA-7B ViT-L/14 8.34B 1.31B 34G 1.5s
@@
LLaMA-3.2-11B-Vision LLaMA-3.1 Vision Adapter 10.67B 10.67B - 23G 3s
MiniCPM-V2.6 Qwen2-7B SigLip-400M 8.10B 8.10B ScienceQA 17G 0.7s
BLIP-2 Flan-T5-XXL ViT-G/14 12.23B 0.11B - 26G 3.5s
InstructBLIP Flan-T5-XXL ViT-G/14 12.31B 0.45B - 16G 0.3s
GLM-4V-9B GLM-4-9B EVA-CLIP-E 13.91B 13.91B ScienceQA 30G 2.4s
InternVL2.5-8B InternViT-300M-448px-V2_5 internlm2_5-7b-chat 8.08B 8.08B ScienceQA 24G 2s
DeepSeek-VL2-tiny MoE-based LLM SigLIP-SO400M 3B 0.57B - 20G 0.2s
DeepSeek-VL2-small MoE-based LLM SigLIP-SO400M 16B 2.4B - 48G 0.25s
DeepSeek-VL2 MoE-based LLM SigLIP-SO400M 27B 4.1B - - -
@@
Qwen2-VL-7B-Instruct Qwen2-7B ViT-G/14 8.29B 8.29B - 18G 0.9s
Qwen2-VL-72B-Instruct Qwen2-72B ViT-G/14 - - - - -
Qwen2.5-VL-3B-Instruct Qwen2.5-3B ViT 3.75B 3.75B - 10G 0.25s
Qwen2.5-VL-7B-Instruct Qwen2.5-7B ViT 8.29B 8.29B - 20G 0.5s